# Positioning-Aided Channel Estimation for Multi-LEO Satellite Downlink Communications

### v1

@@ III-C1 Signal Model
where $\vartheta_{s,u}=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})$ and $\mathbf{z}_{s}[\ell,k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f\mathbf{I}_{N})$ represents the AWGN.
We observe that $\mathbf{d}_{s,u}[\ell,k]$ is directly determined by the UTs’ position feedback. Consequently, instead of estimating the full channel vector, it suffices to estimate only two scalars, i.e., the channel gain $\vartheta_{s,u}$ and the TOA $\tau_{s,u}$. This significantly simplifies the uplink channel estimation task. However, since the position information is based on estimation rather than ground truth, model mismatch arises again. The received signal based on the estimated position information (mismatched model) is expressed as
@@
[…]
@@ III-C2 MCRB and LB Derivation
$\ln{\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto\left\|\mathbf{y}_{s,u}-\bar{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (24)
@@
$\ln{\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto\left\|\mathbf{y}_{s,u}-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (25)
@@
  • Pseudo-True Parameter: According to (6), the pseudo-true parameter, denoted by $\tilde{\bm{\eta}}_{s,u}$, is derived as
@@
    $=\arg\min_{\bm{\eta}_{s,u}}\ \left\|\bar{\mathbf{q}}\left(\bar{\bm{\eta}}_{s,u}\right)-\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (26)
@@
  • MCRB: The generalized FIMs, i.e., $\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}$ and $\mathbf{B}_{\tilde{\bm{\eta}}_{s,u}}$, are obtained by calculating the first- and second-order derivatives of $\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)$ with respect to the elements of $\bm{\eta}_{s,u}$, as described in Appendix B. Based on these, the MCRB is given by
@@
    $\text{LB}\left(\vartheta_{s,u}\right)=\sqrt{\text{Tr}\big{(}\left[\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)\right]_{1:2,1:2}\big{)}},$ (29)
@@
IV Distributed Beamforming for Downlink Communications
The proposed two-timescale PACE enables channel reconstruction between each LEO satellite and all UTs, enabling CSI-based beamforming for downlink communication. To address the limited link budget in single-satellite downlink scenarios, we introduce distributed multi-LEO satellite beamforming where each UT is served by the same satellite cluster. This distributed approach eliminates the need for inter-satellite CSI sharing, as beamforming decisions are made locally at each satellite using only its CSI, rather than requiring consolidated global CSI, which significantly reduces system complexity and latency.
IV-A Signal Model
Without centralized coordination, we manage inter-satellite interference by allocating orthogonal frequency bands to different LEO satellites. The total subcarrier set $\Phi=\{1,2,\ldots,K\}$ is partitioned into $S$ equal subsets $\Phi_{1},\Phi_{2},\ldots,\Phi_{S}$ corresponding to $S$ LEO satellites. For each satellite $s$, its subset is defined as $\Phi_{s}=\{(s-1)\bar{K}+1,(s-1)\bar{K}+1,\ldots,s\bar{K}\}$, where $\bar{K}=K/S$. For the $s$-th LEO satellite transmitting over the $k$-th subcarrier ($k\in\Phi_{s}$), the signal is given by
@@ IV-A Signal Model
where $\mathbf{W}_{s}=[\mathbf{w}_{s,1},\mathbf{w}_{s,2},\ldots,\mathbf{w}_{s,U}]\in\mathbb{C}^{N\times U}$ represents the beamformer [^3] We adopt subcarrier-coherent beamforming, where the beamformer is consistent across subcarriers, with only amplitude adjustments allowed. This approach strikes a balance between performance and computational efficiency, which is critical given the limited onboard processing capabilities of LEO satellites. Although less flexible than subcarrier-specific beamforming, it substantially reduces optimization complexity in systems with a large number of subcarriers. The framework can be extended to subcarrier-specific beamforming, albeit at the expense of increased computational complexity. for the $s$-th LEO satellite, and $\mathbf{s}\left[k\right]\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{U})$ denotes the data streams transmitted to $U$ UTs over the $k$-th subcarrier.
The received signal at the $u$-th UT from the $s$-th LEO satellite over the $k$-th subcarrier can be represented as
$y_{u}\left[k\right]=$ $\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,u}s_{u}\left[k\right]$
$+\underbrace{\sum_{m=1,m\neq u}^{U}\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,m}s_{m}\left[k\right]}_{\text{Inter-user interference}}+z_{u}\left[k\right],$ (32)
where $\mathbf{h}_{s,u}[k]=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})e^{-\jmath 2\pi k\Delta f\tau_{s,u}}\mathbf{a}(\bm{\theta}_{s,u})$ denotes the channel vector incorporating Doppler shift compensation. The term $z_{u}[k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f)$ represents the AWGN. The achievable communication rate for the $u$-th UT, aggregated across all satellites and assigned subcarriers, is given by
$R_{u}=\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}\left[k\right],$ (33)
where $R_{s,u}\left[k\right]$ represents the communication rate between the $s$-th LEO satellite and the $u$-th UT over the $k$-th subcarrier, calculated as
$R_{s,u}\left[k\right]=\log_{2}\left(1+\frac{\left|\mathbf{h}_{s,u}^{\mathsf{T}}\left[k\right]\mathbf{w}_{s,u}\right|^{2}}{I_{s,u}\left[k\right]+N_{0}\Delta f}\right),$ (34)
where $I_{s,u}[k]=\sum_{m=1,m\neq u}^{U}|\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,m}|^{2}$.
@@ IV-B Problem Formulation
$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}R_{u}=\sum_{u=1}^{U}\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}},\;\forall s,$ (35)
@@
Note that, because $\sum_{u=1}^{U}\sum_{s=1}^{S}\sum_{k\in\Phi_{s}}R_{s,u}[k]=\sum_{s=1}^{S}\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}R_{s,u}[k]$, the problem in (IV-B) is inherently distributed. Consequently, a subproblem corresponding to the $s$-th LEO satellite can be expressed as
$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}R_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}}.$ (36)
By solving (IV-B) for each LEO satellite, the original problem in (IV-B) can be effectively addressed.
IV-C Proposed Distributed Beamforming Approach
In the following, we employ the weighted sum mean squared error minimization (WMMSE) optimization framework [36] to solve (IV-B) iteratively. Specifically, by introducing the auxiliary variables $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$, where $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$ serve as the MMSE equalizer and weight, respectively [36], (IV-B) can be equivalently reformulated as
$\mathop{\max}\limits_{\mu_{s,u}[k],\omega_{s,u}[k],\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}\log_{2}\omega_{s,u}[k]-\omega_{s,u}[k]\Psi_{s,u}[k]$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}},$ (37)
where $\Psi_{s,u}[k]=|1-\mu_{s,u}[k]\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}|^{2}+|\mu_{s,u}[k]|^{2}(I_{s,u}[k]+N_{0}\Delta f)$. The optimization problem in (IV-C) is solved by iteratively updating $\mathbf{W}_{s}$, $\omega_{s,u}[k]$, and $\mu_{s,u}[k]$, as detailed below.
IV-C1 Update $\mu_{s,u}[k]$
With $\omega_{s,u}[k]$ and $\mathbf{W}_{s}$ fixed, the optimal $\mu_{s,u}[k]$ is derived by setting $\partial\Psi_{s,u}[k]/\partial\mu_{s,u}[k]=0$, resulting in
$\mu_{s,u}[k]=\frac{\mathbf{h}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}}{I_{s,u}[k]+N_{0}\Delta f}.$ (38)
IV-C2 Update $\omega_{s,u}[k]$
With $\mu_{s,u}[k]$ and $\mathbf{W}_{s}$ fixed, the optimal $\omega_{s,u}[k]$ is obtained by setting $\partial\Psi_{s,u}[k]/\partial\omega_{s,u}[k]=0$, leading to
$\omega_{s,u}[k]=\frac{1}{\Psi_{s,u}[k]}.$ (39)
@@ IV-C3 Update $\mathbf{W}_{s}$
With $\mu_{s,u}[k]$ and $\omega_{s,u}[k]$ fixed, the optimization problem to find the optimal $\mathbf{W}_{s}$ is formulated as
$\mathop{\min}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\sum_{k\in\Phi_{s}}\Psi_{s,u}[k]$
@@
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{\bar{K}}.$ (40)
Here, $\Psi_{s,u}[k]$ can be expressed as the sum of a positive semi-definite quadratic form and a linear term. Consequently, (IV-C3) is a convex quadratically constrained quadratic program (QCQP) problem with respect to $\mathbf{W}_{s}$, which can be efficiently solved via convex optimization toolboxes such as CVX.
The process of solving (IV-B) is outlined in Algorithm 1. The initial beamformer $\mathbf{W}_{s}$ for the $s$-th LEO satellite can be constructed using maximum ratio transmission (MRT), which aligns the beamformer with the reconstructed channels (by excluding the frequency-dependent delay-induced phase rotation). Specifically, we have $\mathbf{W}_{s}=\zeta_{s}[\hat{\vartheta}_{s,1}^{*}\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,1}),\ldots,\hat{\vartheta}_{s,U}^{*}\mathbf{a}^{*}(\hat{\bm{\theta}}_{s,U})]$, where the hat notation indicates estimated values and $\zeta_{s}$ is a normalization factor that ensures the fulfillment of the power budget. Since the WMMSE optimization framework is utilized to solve each subproblem, i.e., (IV-B), its convergence is well-established [36]. The computational complexity of each iteration is primarily determined by solving (IV-C3), which, due to its QCQP nature, is characterized by $\mathcal{O}(N^{3}U^{3})$.
Algorithm 1 Algorithm for Solving (IV-B)
1:Initialize: $\mathbf{W}_{s},\forall s$;
2:for $s=1:S$ do
3:     repeat
4:         Update $\mu_{s,u}[k]$ using (38);
5:         Update $\omega_{s,u}[k]$ using (39);
6:         Update $\mathbf{W}_{s}$ by solving (IV-C3) via CVX;
7:     until the reduction ratio of the objective value falls below a predefined threshold;
8:end for
9:Output: $\mathbf{W}_{s},\forall s$.

### v4

@@ III-B1 Signal Model
where $\mathbf{z}_{s}[\ell,k]\sim\mathcal{CN}(\mathbf{0},N_{0}\Delta f\mathbf{I}_{N})$ represents the AWGN.
We observe that $\mathbf{d}_{s,u}[\ell,k]$ is directly determined by the UTs’ position feedback. Consequently, instead of estimating the full channel vector, it suffices to estimate only two scalars, i.e., the channel gain $\vartheta_{s,u}$ and the TOA $\tau_{s,u}$. [^8] Note that although the TOA $\tau_{s,u}$ has already been estimated during the downlink positioning process, it will nevertheless be estimated during uplink channel estimation as part of standard OFDM processing [53]. This is necessary unless the UT and the LEO satellite are perfectly time-synchronized and the propagation delay remains precisely known, which are hardly satisfied in the highly dynamic LEO satellite environment. This significantly simplifies the uplink channel estimation task. However, since the position information is based on estimation rather than ground truth, model mismatch arises again. The received signal based on the estimated position information (mismatched model) is expressed as
@@
$\tilde{\theta}_{s,u}^{\mathrm{az}}$ $=\text{atan2}\left(\left[\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right]_{2},\left[\tilde{\mathbf{p}}_{u}-\mathbf{p}_{s}\right]_{1}\right),$ (21a)
@@ III-B2 MCRB and LB Derivation
${\xi_{\mathrm{T}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto e^{\left\|\mathbf{y}_{s,u}-\bar{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2}}$ (22)
@@
${\xi_{\mathrm{M}}}\left(\mathbf{y}_{s,u};\bm{\eta}_{s,u}\right)\propto e^{\left\|\mathbf{y}_{s,u}-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2}},$ (23)
@@
  • Pseudo-True Parameter: According to Appendix A, the pseudo-true parameter, denoted by $\tilde{\bm{\eta}}_{s,u}$, is derived as
@@
    $=\arg\min_{\bm{\eta}_{s,u}}\ \left\|\bar{\mathbf{g}}\left(\bar{\bm{\eta}}_{s,u}\right)-\tilde{\mathbf{g}}\left(\bm{\eta}_{s,u}\right)\right\|^{2},$ (24)
@@
  • MCRB: The generalized FIMs, i.e., $\mathbf{A}_{\tilde{\bm{\eta}}_{s,u}}$ and $\mathbf{B}_{\tilde{\bm{\eta}}_{s,u}}$, are obtained by calculating the first- and second-order derivatives of $\tilde{\mathbf{q}}\left(\bm{\eta}_{s,u}\right)$ with respect to the elements of $\bm{\eta}_{s,u}$, as described in Appendix C. Based on these, the MCRB is given by
@@
    $\text{LB}\left(\vartheta_{s,u}\right)=\sqrt{\text{Tr}\big(\left[\text{LBM}\left(\tilde{\bm{\eta}}_{s,u}\right)\right]_{1:2,1:2}\big)}$ (27)
@@
IV Cooperative Beamforming for Downlink Communications
The proposed two-timescale PACE enables channel reconstruction between each LEO satellite and all UTs, enabling CSI-based beamforming for downlink communication. To address the limited link budget in single-satellite downlink scenarios, we introduce multi-LEO satellite cooperative beamforming, where each UT is served by the satellites within the cluster simultaneously. Note that the cooperative beamforming optimization developed in this section is based on a nominal signal model, which treats the channel parameters estimated via PACE as true and assumes perfect delay/Doppler compensation across satellites. In contrast, the communication rate in the simulations is evaluated using the ground-truth channel parameters under imperfect compensation. This separation enables a direct assessment of the effectiveness of PACE by examining the performance achieved by beamformers optimized with nominal (PACE-estimated) parameters. The smaller the performance gap between nominal- and ground-truth-optimized beamformers, the more reliable PACE can be considered. Developing robust beamforming strategies that explicitly account for CSI uncertainty (e.g., [27, 54, 55]) and/or asynchronous interference due to imperfect compensation (e.g., [23]) represents an important but orthogonal research direction, which is left for future work.
IV-A Nominal Signal Model
In the nominal signal model, we assume perfect delay and Doppler compensation across all satellites. Under this assumption, the nominal channel, denoted by $\tilde{\mathbf{h}}_{s,u}=\tilde{\vartheta}_{s,u}\mathbf{a}(\tilde{\bm{\theta}}_{s,u})$, is frequency-flat, where the tilde denotes estimated values. This allows subcarrier-coherent beamforming, where the same beamformer is applied across all subcarriers, thereby reducing optimization complexity. Specifically, for the $s$-th LEO satellite transmitting over the $k$-th subcarrier, the signal is
@@ IV-A Nominal Signal Model
where $\mathbf{W}_{s}=[\mathbf{w}_{s,1},\mathbf{w}_{s,2},\ldots,\mathbf{w}_{s,U}]\in\mathbb{C}^{N\times U}$ represents the beamformer for the $s$-th LEO satellite, and $\mathbf{s}\left[k\right]\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{U})$ denotes the data streams transmitted to the $U$ UTs over the $k$-th subcarrier. The received signal at the $u$-th UT over the $k$-th subcarrier is then
$y_{u}\left[k\right]=\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}s_{u}\left[k\right]+\underbrace{\sum_{m\neq u}^{U}\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}s_{m}\left[k\right]}_{\text{Inter-user interference}}+z_{u}\left[k\right],$ (30)
where $z_{u}[k]\sim\mathcal{CN}(0,N_{0}\Delta f)$ denotes the AWGN.
The nominal achievable communication rate for the $u$-th UT, aggregated across all subcarriers, is given by
$\tilde{R}_{u}=B\log_{2}\left(1+\frac{\left|\sum\limits_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}\right|^{2}}{\sum\limits_{m\neq u}^{U}\left|\sum\limits_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}\right|^{2}+N_{0}\Delta f}\right).$ (31)
@@ IV-B Problem Formulation
$\mathop{\max}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\tilde{R}_{u}$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{K},\;\forall s,$ (32)
@@
Algorithm 1 Proposed Algorithm for Solving (IV-B)
1:Initialize: $\mathbf{W}_{s},\forall s$;
2:repeat
3:  Update $\mu_{u}$ using (34);
4:  Update $\omega_{u}$ using (35);
5:  for $s=1:S$ do
6:   Obtain eigenvalue decomposition of $\mathbf{R}_{s}$;
7:   Determine $\lambda$ by Golden-section search over (53);
8:   Determine $\mathbf{w}_{s}$ via (54) and reshape to obtain $\mathbf{W}_{s}$;
9:  end for
10:until the reduction ratio of the objective value falls below a predefined threshold or a predetermined iteration number is reached;
11:Output: $\mathbf{W}_{s},\forall s$.
IV-C Proposed Cooperative Beamforming Approach
In the following, we employ the WMMSE optimization framework [56] to solve (IV-B) iteratively. Specifically, by introducing the auxiliary variables $\mu_{u}$ and $\omega_{u}$, where $\mu_{u}$ and $\omega_{u}$ serve as the MMSE equalizer and weight, respectively [56], (IV-B) can be equivalently reformulated as
$\mathop{\max}\limits_{\mu_{u},\omega_{u},\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\left(\ln\omega_{u}-\omega_{u}\Psi_{u}\right)$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{K},\;\forall s,$ (33)
where $\Psi_{u}=|1-\mu_{u}\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}|^{2}+|\mu_{u}|^{2}(\sum_{m\neq u}^{U}|\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}|^{2}+N_{0}\Delta f)$. The optimization problem in (IV-C) is solved by iteratively updating $\mathbf{W}_{s}$, $\omega_{u}$, and $\mu_{u}$, as detailed below.
IV-C1 Update $\mu_{u}$
With $\omega_{u}$ and $\mathbf{W}_{s}$ fixed, the optimal $\mu_{u}$ is derived by setting $\partial\Psi_{u}/\partial\mu_{u}=0$, resulting in
$\mu_{u}=\frac{\left(\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,u}\right)^{*}}{\sum_{m=1}^{U}\left|\sum_{s=1}^{S}\tilde{\mathbf{h}}_{s,u}^{\mathsf{T}}\mathbf{w}_{s,m}\right|^{2}+N_{0}\Delta f}.$ (34)
IV-C2 Update $\omega_{u}$
With $\mu_{u}$ and $\mathbf{W}_{s}$ fixed, the optimal $\omega_{u}$ is obtained by setting $\partial\Psi_{u}/\partial\omega_{u}=0$, leading to
$\omega_{u}=\frac{1}{\Psi_{u}}.$ (35)
@@ IV-C3 Update $\mathbf{W}_{s}$
With $\mu_{u}$ and $\omega_{u}$ fixed, the optimization problem to find the optimal $\mathbf{W}_{s}$ is formulated as
$\mathop{\min}\limits_{\mathbf{W}_{s}}\;\;\;$ $\sum_{u=1}^{U}\omega_{u}\Psi_{u}$
${\rm{s.t.}}\;\;\;$ $\sum_{u=1}^{U}\left\|\mathbf{w}_{s,u}\right\|^{2}\leq\frac{P_{s}}{K},\;\forall s.$ (36)
Here, $\Psi_{u}$ can be expressed as the sum of a positive semi-definite quadratic form and a linear term.
We note that (IV-C3) is a convex QCQP with respect to $\mathbf{W}_{s}$, which can be solved using standard toolboxes such as CVX. However, this approach incurs a per-iteration complexity of $\mathcal{O}((SNU)^{3})$, since the beamformers of all satellites are optimized jointly. To better suit computation resource-constrained LEO scenarios, we instead develop a low-complexity method that strategically exploits BCD, strong duality, and matrix inversion properties. [^9] Despite the significant complexity reduction enabled by the proposed scheme, we note that the optimization is still carried out in a centralized manner (e.g., at a cluster-head satellite), which may incur scalability bottlenecks as the network size grows. Designing decentralized strategies that distribute the optimization across participating satellites to improve scalability is an important research direction but lies beyond the scope of this paper. We refer interested readers to our recent work [57] for a scalable distributed beamforming framework over networked LEO satellites. This reduces the dominant complexity of solving (IV-C3) from $\mathcal{O}((SNU)^{3})$ to $\mathcal{O}(SN^{3})$. Further details are provided in Appendix D. The overall procedure for solving (IV-B) is summarized in Algorithm 1, and its convergence follows directly from the BCD structure of the algorithm.
@@
$R_{u}=\Delta f\sum_{k=1}^{K}\log_{2}\!\left(1+\frac{\left|\sum\limits_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,u}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}\!\left[k\right]\mathbf{w}_{s,u}\right|^{2}}{\sum\limits_{m\neq u}^{U}\left|\sum\limits_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,m}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}\!\left[k\right]\mathbf{w}_{s,m}\right|^{2}+N_{0}\Delta f}\right),$ (38)
IV-D Evaluation Model
As stated above, after obtaining the cooperative beamformers by solving the optimization problem under the nominal signal model (with PACE-estimated channel parameters and ideal inter-satellite compensation), we evaluate the achievable communication performance using a more realistic evaluation model that employs the ground-truth channel parameters and explicitly accounts for practical compensation imperfections. This separation is intentional: it enables a direct assessment of the effectiveness of PACE by examining how well beamformers optimized with nominal (PACE-estimated) parameters perform when deployed in a realistic setting.
Following the Doppler-compensation-free assumption adopted in[18, 24], we do not rely on explicit Doppler precompensation at the transmitter in the evaluation model for simplicity. Meanwhile, we account for imperfect delay compensation, which induces residual phase rotations across subcarriers. In addition, motivated by [58, 22, 23], we further incorporate the resulting asynchronous interference due to link-dependent delay mismatch across users in multi-LEO cooperative transmission.
Specifically, after delay precompensation, the received signal of $u$-th UT on $k$-th subcarrier is modeled as [^10] Note that we assume the difference between $\tilde{\tau}_{s,m}$ and $\tau_{s,m}$ is smaller than the cyclic prefix (CP) length.
$y_{u}\!\left[k\right]=\sum_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,u}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,u}\,s_{u}\!\left[k\right]$ (37)
$\quad+\sum_{m\neq u}^{U}\sum_{s=1}^{S}e^{-\jmath 2\pi k\Delta f\left(\tau_{s,u}-\tilde{\tau}_{s,m}\right)}\bar{\mathbf{h}}_{s,u}^{\mathsf{T}}[k]\mathbf{w}_{s,m}\,s_{m}\!\left[k\right]+z_{u}\!\left[k\right],$
where $\bar{\mathbf{h}}_{s,u}[k]=\alpha_{s,u}G(\theta_{s,u}^{\mathrm{el}})\mathbf{a}\!\left(\bm{\theta}_{s,u}\right)$ denotes the (approximately) frequency-flat array response term, while the residual delay mismatch is explicitly captured by the subcarrier-dependent phase rotations in (37).
Accordingly, the achievable rate of $u$-th UT under the evaluation model is given by (38) at the bottom of this page, which makes explicit that asynchronous interference persists due to link-dependent delay mismatch across users, even after applying delay precompensation.
V Numerical Results
V-A Simulation Parameters
We model the Earth as a sphere with a radius of $6400\text{ km}$ and adopt a spherical coordinate system centered at the Earth’s core. The UTs are randomly distributed in a circular service area on the Earth’s surface with a radius of 200 $0\text{ km} [2$8]. The participating LEO satellites are positioned within a corresponding circular area in space, centered along the extension of the line connecting the Earth’s core to the center of the surface service area, at an orbital altitude of $500\text{ km}$. Each satellite travels along an orbital path intersecting this line at a constant speed $\left|\mathbf{v}_{s}\right|=7.6\text{ km/s}$, with velocity vectors $\mathbf{v}_{s}$ tangent to their orbits and UPAs oriented toward Earth’s core. Due to imperfect time-domain and frequency-domain synchronization, the clock bias (respectively, CFO) between each satellite and the UT is randomly set to a value between $8\text{ ns}$ and $12\text{ ns}$ (respectively, between $\Delta f/120$ and $\Delta f/80$), which remains fixed throughout the simulation. To eliminate satellite handover involvement, we consider one large-timescale position-coherent frame containing $100$ small-timescale channel gain-coherent subframes, where the channel gain changes independently from one subframe to another. Throughout this frame, the satellite constellation topology remains constant for simplicity. The downlink positioning and uplink channel estimation consist of $L_{\text{p}}=10000$ and $L_{\text{c}}=1000$ pilot symbols, respectively.
The channel gain from each LEO satellite to each UT, expressed by (3), is independently generated in each channel-gain coherent slot. The random phase $\psi_{s,u}$ is uniformly cooperative in $0,2\pi]$. The components that constitute the path loss $\beta_{s,u}$ are generated according to (4). The free-space path loss is given by $\beta^{\text{FS}}_{s,u}=20\log_{10}(\|\mathbf{p}_{u}-\mathbf{p}_{s}\|)+20\log_{10}(f_{c})-147.55\text{ [dB]}$[41]. The atmospheric absorption $\beta^{\text{AB}}_{s,u}$, which depends on signal frequency and satellite elevation angle, is calculated according to International Telecommunication Union (ITU) recommendations [59] and configured using an off-the-shelf MATLAB script[36]. In the considered LOS-dominant suburban/rural/remote scenarios, the clutter loss $\beta^{\text{CL}}_{s,u}$ and the shadow fading $\beta^{\text{SF}}_{s,u}$ originating from near-surface structures are typically weak[60]. Therefore, we set $\beta^{\text{CL}}_{s,u}=0\text{ dB}$ and $\beta^{\text{SF}}_{s,u}=0\text{ dB}$ for convenience. Additionally, ionospheric scintillation can be ignored for frequencies above $6\text{ GHz}$[41]. For $\beta^{\text{CL}}_{s,u}$ caused by tropospheric scintillation, which is generally difficult to model analytically, we adopt the example suggested in [36] as a reference value. The remaining simulation parameters are listed in Table I.